The energy of a silver atom with a vacancy in K shell is 25.31 keV, in L shell is 3.56 keV and in M shell is 0.530 keV higher than the energy of the atom with no vacancy. Find the frequency of K base α, K base K base β and L base α X-ray of silver.

Grade:11

1 Answers

Aditi Chauhan
9 years ago
E base K = 25.31 KeV, E base L = 3.56 KeV, E base M = 0.530 KeV K base ∝ = E base K – K base L = hv ⇒ v = E base K – E base L/h = 25.31 – 3.56/4.14 * 10^-15 * 10^3 = 5.25 * 10^15 Hz K base β = E base K – K base M = hv ⇒ v = E base K – E base M/h = 23.31 – 0.53/4.14 * 10^-15 * 10^ = 5.985 * 10^18 Hz.